OpenStudy (anonymous):

A rectangular field is 4 times as long as it is wide. If the length is decreased by 10 feet and the width is increased by 2 feet, the perimeter will be 80 feet. Find the dimensions of the original field.

OpenStudy (anonymous):

let original width=x Length=4x again new length=4x-10 new width= x+2 perimeter of rectangle =2(l+w)=2(4x-10+x+2)=80 5x-8=80/2=40 5x=40+8=48 x=48/5=9.6 width=9.6 ft length=4*9.6=38.4 ft